# Two categories of UV-upturn galaxies revealed by semi-analytic models

### v1

@@ 3.5 Joint dependence on $F_{\rm b}$, $F_{\rm y}$, $Z_{\rm*}$ and dust
Figure 9 further explores the joint dependence on $F_{\rm b}$ and stellar metallicity $Z_{\rm*}$. Notably, there is a strong preference for high stellar metallicity in both categories of UV-upturn galaxies: those formed through binary interactions with EHB stars and those resulting from dust attenuation. This finding is consistent with our expectations derived from analyzing SSPs in the YII-ehb model. As discussed in  subsection 3.1, only the most metal-rich SSPs with $Z>0.01$ can evolve into the UV-upturn region after approximately 4 Gyr. In contrast, the latter preference emerges from the complex interactions of star formation activities integrated into semi-analytical and dust attenuation models. In the most extreme case, over 80% of galaxies with $F_{\rm b}>0.7$ and $\log_{10}{(Z_{\rm*}/Z_{\odot})}\sim 0.1$ in panel (d) are classified as UV-upturn galaxies, despite the relatively small absolute number of such metal-rich galaxies. This strong preference for high stellar metallicity accounts for the observed correlation between the UV-upturn phenomenon and metallicities (e.g. Carter et al., 2011; Smith et al., 2012; Akhil et al., 2024). Specifically, we find that the mean stellar metallicities of UV-weak and RSF galaxies in the models are approximately 0.15 to 0.2 dex lower than those of UV-upturn galaxies. This aligns well with the observed metallicity differences between UV-upturn and UV-weak regions, as reported in Jeong et al. (2012).
@@ 4.1 Definition of “UV-upturn”
In fact, the definition or measurement of the UV-upturn phenomenon varies from one paper to another. The relatively stringent photometric definition of UV-upturn galaxies of Yi et al. (2011) has been adopted in some observational (e.g. Jeong et al., 2012; Petty et al., 2013; Boissier et al., 2018) and theoretical (e.g. Hernández-Pérez & Bruzual, 2014) works to physically select UV-upturn galaxies. While in Han et al. (2007) and many other works, the value of the FUV-$r$ color (or other UV-to-optical color) alone was used to quantify the strength of UV excess (e.g. Brown et al., 2003; Donahue et al., 2010; Smith et al., 2012; Phillipps et al., 2020; Ali et al., 2021); likewise, in some other works, the FUV-NUV color alone was used (e.g. Loubser & Sánchez-Blázquez, 2011; Carter et al., 2011). These choices were made sometimes due to the lack of photometric data at certain wavelengths, or because the sample volume might be largely reduced with additional color criteria. Moreover, observational systematic uncertainties further constrain the power of rather stringent criteria. For example, when the UV-upturn criteria of Yi et al. (2011) were applied to the Coma cluster, Yi et al. (2011) found that none of the 30 brightest elliptical galaxies falls into the UV-upturn region. In contrast, Smith et al. (2012) reported that 14 out of 150 quenched galaxies in the Coma cluster meet the criteria, and this discrepancy was likely due to their different applications of k-correction. Furthermore, in some spectroscopic studies of ETGs at high redshift (e.g. Lonoce et al., 2020), line indices were used in SED fitting to ascertain whether a galaxy favors an old stellar population with an upturn contribution, in which case it would be defined as a UV-upturn galaxy. Le Cras et al. (2016) found that the fraction of UV-upturn galaxies defined in this way was approximately seven times that reported in Yi et al. (2011), as some UV-weak galaxies might also select an upturn SSP if only the line indices’ strength were considered.
@@ 4.4 Limitations and outlook
Utilizing the YII-ehb EPS model, our semi-analytic model gabe roughly reproduced the observed bimodal color distributions of elliptical galaxies. However, it yields a much narrower FUV-NUV range and an inaccurately positioned star-forming sequence, as shown in Figure 3. Achieving a fully reproduction of the observed color distributions of galaxies remains a difficult challenge in semi-analytic models. Firstly, galaxy colors are generally not used as a model calibrator. Instead, the stellar mass function is the predominant observational result employed for calibration in semi-analytic models (e.g. Guo et al., 2011). Furthermore, scaling relations, such as the supermassive black hole to stellar bulge mass relation, the galaxy size-mass relation, and the metallicity-mass relation, are also ensured during the calibration process. In contrast to these carefully treated galaxy properties, galaxy colors are predictions of semi-analytic models, rather than calibrations. Consequently, semi-analytic models may not perform their best in predicting galaxy colors within the feasible parameter space. Secondly, galaxy colors are highly dependent on the SFHs of galaxies, which are challenging to reproduce self-consistently across various types of galaxies and across different redshifts. Unlike integral galaxy properties such as stellar mass, the star formation rate is differential and sensitive to almost all galactic physical processes, encompassing reionization, hot gas cooling, supernova feedback, galaxy mergers, etc., especially tidal and ram-pressure stripping in satellite galaxies and AGN feedback in massive galaxies. Lastly, numerous physical processes that might influence galaxy colors have not been exquisitely considered in semi-analytic models. As discussed in previous sections, incorporating variations of IMFs, dust attenuation, and binary fractions can significantly alter the distribution of model galaxies on the color-color diagram. The narrower distributions of model galaxies in comparison to the observed one actually reflect a huge amount of physical processes and their variations in the real Universe which are missed in current semi-analytic models.
@@ 5 Conclusions
UV-upturn galaxies are elliptical galaxies (or central bulges in disk galaxies) with abnormally excessive flux in the FUV band, which is generally ascribed to the formation of EHB stars in old stellar populations. By using the semi-analytic model of galaxy formation and evolution gabe (Jiang et al., 2019), we embed the binary star formation model of EHB stars (Han et al., 2002, 2003, 2007; Zhang et al., 2020) within a framework of cosmological evolution to introduce variances of SFHs and explore the formation and evolution of UV-upturn galaxies in a cosmological context. We have gone through all physical mechanisms that are possibly related to the UV-upturn phenomenon, including stellar evolution (EPS models), initial mass function (IMF) and dust attenuation, as well as the age, metallicity and binary fractions of stellar populations in a galaxy, attempting to figure out which processes play a more important role in forming/identifying a UV-upturn galaxy.
@@
  • Having examined vaious EPS models including binary and single star versions of yunnan-ii, bpass, as well as galaxev, we find that only the yunnan-ii model which has considered the formation of EHB stars through binary interactions (YII-ehb) could provide enough FUV flux in old stellar populations and reproduce a relatively large fraction of UV-upturn galaxies.

### v2

@@ 3.5 Joint dependence on $F_{\rm b}$, $F_{\rm y}$, $Z_{\rm*}$ and dust
Figure 9 further explores the joint dependence on $F_{\rm b}$ and stellar metallicity $Z_{\rm*}$. Notably, there is a strong preference for high stellar metallicity in both categories of UV-upturn galaxies: those formed through binary interactions with EHB stars and those resulting from dust attenuation. This finding is consistent with our expectations derived from analyzing SSPs in the YII-ehb model. As discussed in subsection 3.1, only the most metal-rich SSPs with $Z>0.01$ can evolve into the UV-upturn region after approximately 4 Gyr. In contrast, the latter preference emerges from the complex interactions of star formation activities integrated into semi-analytical and dust attenuation models. In the most extreme case, over 80% of galaxies with $F_{\rm b}>0.7$ and $\log_{10}{(Z_{\rm*}/Z_{\odot})}\sim 0.1$ in panel (d) are classified as UV-upturn galaxies, despite the relatively small absolute number of such metal-rich galaxies. This strong preference for high stellar metallicity accounts for the observed correlation between the UV-upturn phenomenon and metallicities (e.g. Carter et al., 2011; Smith et al., 2012; Akhil et al., 2024). Specifically, we find that the mean stellar metallicities of UV-weak and RSF galaxies in the models are approximately 0.15 to 0.2 dex lower than those of UV-upturn galaxies. This aligns well with the observed metallicity differences between UV-upturn and UV-weak regions, as reported in Jeong et al. (2012). We will further discuss on the implications of the metallicity dependence of the UV-upturn phenomenon in subsection 4.2.
@@ 4.1 Influence of the definition of “UV-upturn”
As pointed out in subsection 2.2, the definition or measurement of the UV-upturn phenomenon varies from one paper to another. The relatively stringent photometric definition of UV-upturn galaxies of Yi et al. (2011) has been adopted in some observational (e.g. Jeong et al., 2012; Petty et al., 2013; Boissier et al., 2018) and theoretical (e.g. Hernández-Pérez & Bruzual, 2014) works to physically select UV-upturn galaxies. While in Han et al. (2007) and many other works, the value of the FUV-$r$ color (or other UV-to-optical color) alone was used to quantify the strength of UV excess (e.g. Brown et al., 2003; Donahue et al., 2010; Smith et al., 2012; Phillipps et al., 2020; Ali et al., 2021); likewise, in some other works, the FUV-NUV color alone was used (e.g. Loubser & Sánchez-Blázquez, 2011; Carter et al., 2011). These choices were made sometimes due to the lack of photometric data at certain wavelengths, or because the sample volume might be largely reduced with additional color criteria. Moreover, observational systematic uncertainties further constrain the power of rather stringent criteria. For example, when the UV-upturn criteria of Yi et al. (2011) were applied to the Coma cluster, Yi et al. (2011) found that none of the 30 brightest elliptical galaxies falls into the UV-upturn region. In contrast, Smith et al. (2012) reported that 14 out of 150 quenched galaxies in the Coma cluster meet the criteria, and this discrepancy was likely due to their different applications of k-correction. Furthermore, in some spectroscopic studies of ETGs at high redshift (e.g. Lonoce et al., 2020), line indices were used in SED fitting to ascertain whether a galaxy favors an old stellar population with an upturn contribution, in which case it would be defined as a UV-upturn galaxy. Le Cras et al. (2016) found that the fraction of UV-upturn galaxies defined in this way was approximately seven times that reported in Yi et al. (2011), as some UV-weak galaxies might also select an upturn SSP if only the line indices’ strength were considered.
@@ 4.4 Limitations and outlook
Utilizing the YII-ehb EPS model, our semi-analytic model gabe roughly reproduced the observed bimodal color distributions of elliptical galaxies. However, it yields a much narrower FUV-NUV range and an inaccurately positioned star-forming sequence, as shown in Figure 3. Achieving a fully reproduction of the observed color distributions of galaxies remains a difficult challenge in semi-analytic models. Firstly, galaxy colors are generally not used as a model calibrator. Instead, the stellar mass function is the predominant observational result employed for calibration in semi-analytic models (e.g. Guo et al., 2011). Furthermore, scaling relations, such as the supermassive black hole to stellar bulge mass relation, the galaxy size-mass relation, and the metallicity-mass relation, are also ensured during the calibration process. In contrast to these carefully treated galaxy properties, galaxy colors are predictions of semi-analytic models, rather than calibrations. Consequently, semi-analytic models may not perform their best in predicting galaxy colors within the feasible parameter space. Secondly, galaxy colors are highly dependent on the SFHs of galaxies, which are challenging to reproduce self-consistently across various types of galaxies and across different redshifts. Unlike integral galaxy properties such as stellar mass, the star formation rate is differential and sensitive to almost all galactic physical processes, encompassing reionization, hot gas cooling, supernova feedback, galaxy mergers, etc., especially tidal and ram-pressure stripping in satellite galaxies and AGN feedback in massive galaxies. Lastly, numerous physical processes that might influence galaxy colors have not been exquisitely considered in semi-analytic models. As discussed in previous sections, incorporating variations of IMFs, dust attenuation, and binary fractions can significantly alter the distribution of model galaxies on the color-color diagram. The narrower distributions of model galaxies in comparison to the observed one actually reflect such physical processes and their variations in the real Universe which are missed in current semi-analytic models.
@@ 5 Conclusions
UV-upturn galaxies are elliptical galaxies (or central bulges in disk galaxies) with abnormally excessive flux in the FUV band. Although the origin of such UV-upturn phenomenon is still under debate, it is generally ascribed to the formation of EHB stars and their hot descendants in old stellar populations (See section 1 for a brief summary of the literature). By using the semi-analytic model of galaxy formation and evolution gabe (Jiang et al., 2019), we embed the binary star formation model of EHB stars (Han et al., 2002, 2003, 2007; Zhang et al., 2020) within a framework of cosmological evolution to introduce variances of SFHs and explore the formation and evolution of UV-upturn galaxies in a cosmological context. We have gone through all physical mechanisms that are possibly related to the UV-upturn phenomenon, including stellar evolution (EPS models), initial mass function (IMF) and dust attenuation, as well as the age, metallicity and binary fractions of stellar populations in a galaxy, attempting to figure out which processes play a more important role in forming/identifying a UV-upturn galaxy.
@@
  • Having examined vaious EPS models including binary and single star versions of yunnan-ii, bpass, as well as galaxev, we find that only the yunnan-ii model which has considered the formation of EHB stars through binary interactions (YII-ehb) could provide enough FUV flux in old stellar populations and reproduce a relatively large fraction of UV-upturn galaxies. However, all these models are actually not capable of reproducing the entire FUV-NUV range of colors observed in the NSA sample of elliptical galaxies (Figure 3).
@@ Acknowledgements
Appendix A Redshift evolution of UV-optical colors
A.1 Dependence on dust attenuation model
Figure 10: Dust attenuation curves as parameterized in LABEL:tab:dust_models. Tan solid line is our fiducial model (Mathis et al., 1983); tan dashed and dot-dashed line are the contributions of graphite and silicate respectively. Red and blue solid lines are dust curves with different UV bump amplitude $E_{\rm b}$ and slope $\delta$ respectively. Vertical dashed lines mark the effective wavelengths of the FUV (1528 Å), NUV (2271 Å, Morrissey et al. 2005) and r band (6230 Å, York et al. 2000).
Figure 11: Redshift evolution of the rest-frame NUV-$r$ (left) and rest-frame FUV-$r$ colors (right) of modeled quenched galaxies (sSFR $<10^{-12}\,{\rm yr}^{-1}$) run with the YII-ehb EPS model and various dust attenuation models as parameterized in LABEL:tab:dust_models. Symbols with error bars show the observational measurements, repeated from Figure 4. Lower panels show the differences between different dust models and the fiducial M83 model. The tan horizontal line indicates zero.
Figure 12: Redshift evolution of the rest-frame NUV-$r$ (left) and rest-frame FUV-$r$ colors (right) of quenched model galaxies (sSFR $<10^{-12}\,{\rm yr}^{-1}$), run with YII-sin, BPASS-sin, and BPASS-bin EPS model with various IMFs. Symbols with error bars show the observational measurements, repeated from Figure 4. Lower panels show the differences between different IMFs and the one with $\Gamma_{\rm bi}=$ 1.3 (for YII-sin) or 1.35 (for BPASS-sin and BPASS-bin).
Dust attenuation curves as parameterized in LABEL:tab:dust_models are shown in Figure 10. As described in subsection 3.4, in addition to our fiducial curve of Mathis et al. (1983), we consider two sets of models of dust attenuation curves with varying parameters $\delta$ and $E_{\rm b}$. The first set (E130, E035, E000) is aimed to isolate the effects of UV bump strength, with a fixed slope of $\delta=0$ but three different values of $E_{\rm b}$, as shown in the left panel. The second set of models (Dm06, D00, Dp03) are assumed to follow the observed relation of $E_{\rm b}=-1.9\delta+0.85$ from Kriek & Conroy (2013) to explore the effects of varying slopes with a relatively weak bump, as shown in the right panel. Vertical dashed lines in Figure 10 mark the effective wavelengths of the FUV, NUV, and $r$ band respectively. It is evident that the UV bump strength mainly influences the extinction in the NUV band, applying little effect in the FUV band. With a stronger UV bump, the NUV-$r$ color becomes redder while the FUV-NUV color becomes bluer, which agrees with the behavior of UV-upturndust galaxies in the E130 model of Figure 7. In contrast, with a steeper slope, both the FUV-$r$ and NUV-$r$ colors become redder, and the extinction in the FUV band is even stronger than in the NUV band.
Figure 11 shows the redshift evolution of the rest-frame NUV-$r$ (left) and rest-frame FUV-$r$ colors (right) of modeled quenched galaxies (sSFR $<10^{-12}\,{\rm yr}^{-1}$) run with the YII-ehb EPS model and various dust attenuation models as parameterized in LABEL:tab:dust_models. The differences between different dust models and the fiducial M83 model are in the lower panels. We find that different dust models merely have a tiny influence on the UV-to-optical colors of these quenched galaxies, with $\Delta$(NUV-$r$) $\lesssim 0.10$ and $\Delta$(FUV-$r$) $\lesssim 0.15$, as these quenched galaxies are naturally devoid of dust. The differences are much smaller than the observational uncertainties, as indicated by the data points with error bars, hence negligible.
Although the influence is slight, the overall trend agrees with that of Figure 10. Dust attenuation model with a stronger UV bump (E130) results $\sim 0.05$ redder NUV-$r$ colors, while it has almost no influence on the FUV-$r$ colors. In contrast, the one with a steeper slope (Dm06) results both redder NUV-$r$ and FUV-$r$ colors, and the impact on the FUV-$r$ colors is larger.
A.2 Dependence on initial mass function
Similar to Figure 11, Figure 12 shows the redshift evolution of the rest-frame NUV-$r$ (left panel) and FUV-$r$ (right panel) colors for modeled quenched galaxies, run with the YII-sin, BPASS-sin, and BPASS-bin EPS model with various IMFs, as indicated. Note that YII-ehb EPS models are not included in this analysis for the reason explained in subsection 3.3. As can be seen from Figure 12, all the models considered predict redder UV-to-optical colors than the observation. This is consistent with the conclusion in the main part of the paper that only the YII-ehb model which has considered the formation of EHB stars through binary interactions could provide enough FUV flux in old stellar populations. Overall, it is clear that the variations in the colors between models of different IMFs are relatively small, less than $\sim 0.1$ magnitude in all cases except the YII-sin model with $\Gamma=0.4$, which predicts even redder FUV-$r$ colors than the same models but with steeper IMFs. These results reinforce the conclusion that variations in the high-mass end slope of IMFs have a negligible effect on the UV-upturn phenomenon.